OpenStudy (anonymous):

Last question of the night, can someone PLEASE tell me if I am on the right track with this problem?? A certain mountain has an elevation of 19,338 feet. In 1921, the glacier on this peak covered 7 acres. By 2001 this glacier had melted to only 1 acre. A) assume that this glacier melted at a constant rate each year. B) Use youir answer for part A to write a linear equation that give the acreage A of this glaicer t years past 1921. a-the yearly rate of change os __________ acres/year b-what is the equation that gives the acreage of the glacier t years after 1921?

OpenStudy (anonymous):

So far I have: m = (y-y)/ (x-x) = 7-1 (1921, 7) & (2001, 1) Am i on the right track?